Circular No. 6224

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP)

Continuation to the ephemeris on IAUC 6202:

1995TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Sept. 30 18^{h}16.^{m}16 −29^{o}20.^{′}5 6.536 6.562 87.^{o}1 8.^{o}8 10.2
Oct. 10 18 16.66 −28 52.7 6.618 6.474 77.4 8.7 10.2
20 18 18.26 −28 25.4 6.693 6.386 67.9 8.3 10.2
30 18 20.87 −27 58.6 6.758 6.297 58.6 7.7 10.1
Nov. 9 18 24.37 −27 32.1 6.807 6.208 49.4 7.0 10.1
19 18 28.67 −27 05.8 6.838 6.118 40.3 6.0 10.0
29 18 33.64 −26 39.4 6.848 6.027 31.4 4.9 10.0

This ephemeris continuation is from new orbital elements computed from 597 observations extending to Sept. 3 and issued today on MPC 25623. The near-perihelic (T = 1997 Mar. 31.96) ephemeris below revises that on IAUC 6202. Although the computation depends heavily on R. H. Mc-Naught’s measurement of a presumed image of the comet in 1993 (IAUC 6198), an orbit computed independently from the 1995 observations alone does now confirm that the comet can not be ‘new’ in the Oort sense. There is thus some justification for the use of an r^{−4} law for the computation of the total magnitude m_{1}, although the comet could easily be two magnitudes fainter (or, conceivably, brighter) than indicated. R. M. West, European Southern Observatory, reports that 1-m Schmidt plates taken on 1993 May 26 and 1994 Aug. 4 just miss the region where the comet would have been. It therefore seems unlikely that other pre-1995 images of the comet exist.

1997 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Feb. 1 19^{h}40.^{m}62 +15^{o}32.^{′}1 2.002 1.366 37.^{o}9 26.^{o}3  0.9
11 20 08.00 +20 54.0 1.805 1.250 41.5 31.5  0.2
21 20 43.91 +27 26.5 1.617 1.142 44.3 37.2 −0.4
Mar. 3 21 33.91 +34 53.2 1.458 1.049 46.0 42.8 −1.0
13 22 45.56 +41 54.5 1.350 0.975 46.1 47.3 −1.5
23 0 20.05 +45 42.1 1.315 0.928 44.8 49.1 −1.7
Apr. 2 1 57.42 +44 06.3 1.359 0.915 42.3 47.4 −1.7
12 3 13.80 +38 32.1 1.469 0.935 39.0 42.5 −1.5
22 4 06.90 +31 45.6 1.620 0.988 35.3 36.0 −1.0
May 2 4 44.35 +25 17.8 1.789 1.067 31.4 29.5 −0.5
12 5 12.53 +19 33.7 1.960 1.164 27.6 23.8  0.1
22 5 35.20 +14 31.2 2.123 1.273 24.5 19.2  0.7
June 1 5 54.46 +10 01.4 2.274 1.390 22.3 16.1  1.2

1995 September 9 Brian G. Marsden